OpenStudy (anonymous):

determine the exact value of cos465°

OpenStudy (anonymous):

\[a) \frac{ \sqrt{3}+\sqrt{2} }{ 2 }\]\[b) \frac{ \sqrt{2}+\sqrt{3}\sqrt{2}}{ 4 }\]\[c) \frac{ \sqrt{2}-\sqrt{3}\sqrt{2} }{ 4 }\]\[d) \frac{ 2+\sqrt{2}-\sqrt{3}\sqrt{2} }{ 4 }\]

OpenStudy (dumbcow):

\[\cos 465 = \cos 105 = \cos(60+45)\] angle sum identity: \[\cos(a+b) = \cos a \cos b - \sin a \sin b\]

OpenStudy (dumbcow):

hopefully that helps...you can get the values from a unit circle only tricky part may be in simplifying the fraction

OpenStudy (anonymous):

lol im trying to simplify it now xD

OpenStudy (anonymous):

i got c ... is that right? :S @dumbcow

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

really? YAY! thanks!

OpenStudy (dumbcow):

yw